# ICNN-enhanced 2SP: Leveraging input convex neural networks for solving two-stage stochastic programming

### v1

@@ Abstract
Two-stage stochastic programming (2SP) offers a basic framework for modelling decision-making under uncertainty, yet scalability remains a challenge due to the computational complexity of recourse function evaluation. Existing learning-based methods like Neural Two-Stage Stochastic Programming (Neur2SP) employ neural networks (NNs) as recourse function surrogates but rely on computationally intensive mixed-integer programming (MIP) formulations. We propose ICNN-enhanced 2SP, a method that leverages Input Convex Neural Networks (ICNNs) to exploit linear programming (LP) representability in convex 2SP problems. By architecturally enforcing convexity and enabling exact inference through LP, our approach eliminates the need for integer variables inherent to the conventional MIP-based formulation while retaining an exact embedding of the ICNN surrogate within the 2SP framework. This results in a more computationally efficient alternative that maintains solution quality. Comprehensive experiments reveal that ICNNs incur only marginally longer training times while achieving validation accuracy on par with their MIP-based counterparts. Across benchmark problems, ICNN-enhanced 2SP often exhibits considerably faster solution times than the MIP-based formulations while preserving solution quality, with these advantages becoming significantly more pronounced as problem scale increases. For the most challenging instances, the method achieves speedups of up to 100$\times$ and solution quality superior to MIP-based formulations.
@@ 1 Introduction
Two-stage stochastic programming (2SP) is a cornerstone modelling framework for decision-making under uncertainty, enabling optimal resource allocation in the face of unpredictable future events. In this, initial (first-stage) decisions are made before uncertainty is resolved, followed by adjustment (recourse, or second-stage) decisions once the random variables have unfolded. Industries such as power (Römisch and Vigerske, 2010) and petrochemical sectors (Khor et al., 2008) rely on 2SP to incorporate uncertainty in their optimisation modelling, utilising it as a standard tool in planning activities.
@@
Sample Average Approximation (SAA) (Shapiro and Homem-de Mello, 1998) is a foundational technique for 2SP, in which the expected value of the second-stage problem is replaced with a finite-scenario approximation. Though computationally tractable, SAA faces a fundamental trade-off: insufficient scenarios may fail to accurately represent the random phenomena, while larger samples increase computational demands. To alleviate this limitation, scenario reduction techniques (Heitsch and Römisch, 2003) are often applied, but these risk omitting critical scenarios, undermining solution robustness. These shortcomings have prompted the adoption of specialised decomposition approaches, including Benders-based (Sherali and Fraticelli, 2002) and Lagrangian decomposition-based methods (Carøe and Schultz, 1999), which decompose the problem into smaller, more manageable subproblems. By exploiting problem structure (e.g., separability and duality), these methods mitigate SAA’s scalability challenges and reduce dependency on scenario sampling (Küçükyavuz and Sen, 2017). However, their efficacy remains contingent on problem-specific properties, such as linearity or convexity, and often requires meticulous algorithmic tuning. Despite developments in decomposition methods (Varga et al., 2024), these limitations persist, driving exploration of alternative approaches that leverage machine learning (ML).
@@
Recent advances in learning-based methods for 2SP have explored surrogate models to approximate complex value functions. Dai et al. (2021) proposed a neural network (NN) to map problem instances to piecewise-linear value functions within low-dimensional latent spaces. Larsen et al. (2024) replace later-stage Benders subproblems with ML predictions to accelerate decomposition. Yilmaz and Büyüktahtakın (2024) decompose 2SP into multi-agent structures trained on sampled scenarios. While these approaches demonstrate the advantages of surrogate modelling, they are relatively less generalisable, constrained by specific problem components or reliance on bespoke architectures. In contrast, Neur2SP (Neural Two-Stage Stochastic Programming) (Dumouchelle et al., 2022) presents a general framework by employing NNs to approximate the expected recourse function, motivated by the embedding of a trained NN surrogate (Fischetti and Jo, 2018). However, Neur2SP’s generality comes at a cost: embedding NNs with rectified linear unit (ReLU) activations into optimisation models requires mixed-integer programming (MIP) formulations that scale with the network size, leading to computational intractability for deeper or wider architectures. This tension between generality and efficiency underscores the need for architectures that balance expressivity with tractability.
@@
Input Convex Neural Networks (ICNNs) (Amos et al., 2017) have emerged as a promising alternative for surrogate modelling due to their ability to fit general convex functions while retaining the flexibility and expressivity of NNs. Unlike typical feed-forward networks, ICNNs are architecturally constrained to ensure convexity with respect to their inputs, enabling exact inference through linear programming (LP) formulations (Duchesne et al., 2021). This property renders them particularly advantageous for embedding within large-scale optimisation frameworks. Although ICNNs cannot model non-convex recourse functions, a limitation compared to ReLU NNs, many real-world 2SP problems in inventory management, energy dispatch, and portfolio optimisation inherently exhibit convex structure (Agrawal and Jia, 2019; Zhong et al., 2021; Luxenberg and Boyd, 2024). Despite their promise, prior research has predominantly focused on applying ICNNs for function approximation and system modelling, such as learning Optimal Power Flow (OPF) value functions (Chen et al., 2020; Rosemberg et al., 2024) or representing system dynamics in optimal control (Yang and Bequette, 2021; Bünning et al., 2021; Jiang et al., 2022; Wang et al., 2024; Omidi et al., 2024). However, their potential in surrogate-based optimisation, specifically for approximating convex recourse functions in 2SP, remains underexplored. As such, our work bridges this gap by unlocking efficient and scalable solutions for 2SP, leveraging the inherent convexity and computational tractability of LP.
@@
In particular, we present several interconnected contributions within a surrogate-based optimisation framework by integrating ICNNs into the 2SP paradigm, as we illustrate in Figure 1. First, we exploit the LP representability of ICNNs to eliminate integer variables inherent to conventional MIP-based embeddings. Second, we propose the first integration of ICNNs in 2SP, training them to approximate the expected value function while preserving convexity. Third, extensive numerical experiments show that our method achieves solution quality comparable to Neur2SP (Dumouchelle et al., 2022) while reducing computational time, especially for large-scale instances, enabling application to time-critical decision problems.
@@ 2.1 Two-stage stochastic programming
which evaluates to the optimal second-stage cost under realisation $\xi$ given the first-stage decision $x$. We assume relatively complete recourse, ensuring $Q$ is feasible for any $x$ satisfying (2.1) and (2.1).
@@ 2.2 ICNN architectures
The surrogate function $\hat{f}$ is guaranteed to be convex in $z_{0}$ provided that $W_{1:K-1}$ are non-negative and the activation function $\sigma(\cdot)$ (ReLU in our case) is convex and non-decreasing (Amos et al., 2017). This follows standard results relating to convexity-preserving operations applied to convex functions: (i) non-negative sums of convex functions are convex; (ii) the composition of a convex function with a convex non-decreasing function retains convexity. Another relevant feature of ICNNs is the presence skip connections parameterised by $S_{k}$, which directly propagates the input $z_{0}$ to deeper layers. This design aims to mitigate the restrictive effects of the non-negativity constraints imposed on $W_{k}$. Notably, these skip connections preserve convexity as the sum of a convex function (from previous layers) and a linear function (from the skip connections) remains convex.
@@ 2.3 Exact inference in ICNNs
In this formulation, the non-linearity of ReLU activations and layer-wise equality constraints are transformed into inequalities via an epigraph representation, exploiting the fact that ReLU is a piecewise-linear function. One may note that the minimisation objective enforces that at least one constraint in (2.3) or (2.3) becomes active at the optimal solution, recovering the original network’s feedforward behaviour. This transformation provides a principled framework for embedding a trained ICNN surrogate into larger optimisation problems, as it offers not only an exact representation of the NN but also preserves computational tractability inherent to LP.
@@ 3 Convexity analysis
This section establishes the theoretical foundation for using ICNNs as surrogates for the second-stage cost $\mathcal{Q}(x)$ in 2SP. Since ICNNs inherently preserve convexity, their effectiveness hinges on whether $\mathcal{Q}(x)$ is convex. As $\mathcal{Q}(x)$ is evaluated by solving the second-stage problem $Q(x,\xi)$ for each scenario, it becomes a convex function of $x$ only if the second-stage problem itself is convex. A convex problem can take two equivalent forms: minimising a convex function or maximising a concave function - as these are mathematically equivalent through the negation of the objective - over a convex feasible region. We analyse three variants of 2SP problems to demonstrate that $\mathcal{Q}(x)$ retains its convexity under broad conditions, even when discrete variables are present in the second-stage problem.
Case 1: continuous first- and second-stage variables. When $x\in\mathbb{R}^{n}_{+}$ and $y_{s}\in\mathbb{R}^{n_{y}}_{+}$, $\forall s\in S$, the second-stage problem $Q(x,\xi_{s})$ is an LP. By LP duality, we have that
@@
Each term $-\pi^{\intercal}_{k}T_{s}x+\pi^{\intercal}_{k}h_{s}$ is affine in $x$, so $Q(x,\xi_{s})$ is the pointwise maximum of affine functions. As affine functions are convex and the pointwise maximum preserves convexity (Boyd and Vandenberghe, 2004, Section 3.2.3), $Q(x,\xi_{s})$ is convex in $x$. The expected second-stage cost $\mathbb{E}_{\xi}[Q(x,\xi)]=\sum_{s\in S}p_{s}Q(x,\xi_{s})$ is a convex combination of convex functions $Q(x,\xi_{s})$, since $0\leq p_{s}\leq 1$, $\forall s\in S$, and $\sum_{s\in S}p_{s}=1$. Thus, $\mathcal{Q}(x)$ is convex in $x$.
Case 2: mixed-integer first- and continuous second-stage variables. Consider $x\in\mathcal{X}$ where $\mathcal{X}=\{x\in\mathbb{Z}^{n}_{+}\mid Ax=b\}$ and $y_{s}\in\mathbb{R}^{n_{y}}_{+}$. Let us define $\mathcal{X}_{\text{relax}}=\{x\in\mathbb{R}^{n}_{+}\mid Ax=b\}$ as the continuous relaxation of $\mathcal{X}$, and $\textbf{conv}(\mathcal{X})$ as the convex hull of $\mathcal{X}$. From Case 1, we established that $\mathcal{Q}(x)$ is convex over $\mathcal{X}_{\text{relax}}$. Since $\textbf{conv}(\mathcal{X})\subset\mathcal{X}_{\text{relax}}$, the function $\mathcal{Q}(x)$ is also convex over $\textbf{conv}(\mathcal{X})$. Thus, there must be $\lambda$ values with $\lambda\in[0,1]$ for which all points in $\mathcal{X}$ can be represented as $x_{\lambda}=\lambda x_{1}+(1-\lambda)x_{2}$ for $x_{1},x_{2}\in\textbf{conv}(\mathcal{X})$. As we have established that $\mathcal{Q}(x)$ is convex over $\textbf{conv}(\mathcal{X})$, the convexity property $\mathcal{Q}(x_{\lambda})\leq\lambda\mathcal{Q}(x_{1})+(1-\lambda)\mathcal{Q}(x_{2})$ must hold for these points $x_{\lambda}$ and their respective $\lambda$ values. Thus, as $x_{\lambda}$ collectively form $\mathcal{X}$, $\mathcal{Q}(x)$ is convex over $\mathcal{X}$, though not continuous due to the discrete nature of $\mathcal{X}$. We highlight that, when $x$ is mixed-integer, the convexity of $\mathcal{Q}(x)$ trivially applies by the same reasoning.
Case 3: mixed-integer first- and second-stage variables. Following Case 2, we first demonstrate that $Q(x,\xi_{s})$ remains convex in $x$ when $y_{s}\in\mathbb{Z}^{n_{y}}_{+}$. For a fixed $x\in\mathcal{X}$, we define the feasible region of the second-stage problem as:
$F_{s}(x)=\{y_{s}\in\mathbb{Z}^{n_{y}}_{+}\mid W_{s}y_{s}=h_{s}-T_{s}x\}.$ (5)
Let $\textbf{conv}(F_{s}(x))$ be the convex hull of this set. Since the optima lie at extreme points of the feasible set in LP; and the extreme points of $\textbf{conv}(F_{s}(x))$ are precisely the integer points in $F_{s}(x)$, we have:
$Q(x,\xi_{s})=\min_{y_{s}\in F_{s}(x)}q_{s}^{\intercal}y_{s}=\min_{y_{s}\in\textbf{conv}(F_{s}(x))}q_{s}^{\intercal}y_{s}.$ (6)
With this equivalence shown, we can apply LP duality as in Case 1. Following the same reasoning from Case 1, we conclude $Q(x,\xi_{s})$ is the pointwise maximum of affine functions in $x$, and thus convex. Consequently, $\mathcal{Q}(x)=\sum_{s\in S}p_{s}Q(x,\xi_{s})$ remains convex in $x$. For mixed-integer second-stage variables, the resulting problem combines the structures analysed in Cases 1 and the integer second-stage variables case. Thus, our convexity analysis extends naturally to mixed-integer recourse.
Remark.
@@
The convexity of $\mathcal{Q}(x)$ extends to second-stage problems with nonlinear convex constraints, provided that appropriate constraint qualification (CQ) conditions are satisfied, such as Slater’s CQ condition (Boyd and Vandenberghe, 2004, Section 5.2.3), since strong duality still applies under CQ. In this more general setting, the dual formulation transforms $Q(x,\xi_{s})$ into the pointwise maximum of convex functions in $x$. Following similar reasoning as in Case 1, $\mathcal{Q}(x)$ preserves its convexity. Furthermore, if the first-stage problem includes a nonlinear convex objective, along with nonlinear convex inequality constraints and affine equality constraints, the recourse function remains convex in $x$ (Birge and Louveaux, 2011, Section 3.5).

### v2

@@ Abstract
[…]
keywords:
stochastic optimisation , surrogate modelling , neural networks
[…]
@@ 1 Introduction
Two-stage stochastic programming (2SP) is a cornerstone modelling framework for decision-making under uncertainty, enabling optimal resource allocation in the face of unpredictable future events. In this, initial (first-stage) decisions are made before uncertainty is resolved, followed by adjustment (recourse, or second-stage) decisions once the random variables have unfolded. Industries such as power [romisch_recent_2010] and petrochemical sectors [khor_two-stage_2008] rely on 2SP to incorporate uncertainty in their optimisation modelling, utilising it as a standard tool in planning activities.
@@
Sample Average Approximation (SAA) [shapiro_simulation-based_1998] is a foundational technique for 2SP, in which the expected value of the second-stage problem is replaced with a finite-scenario approximation. Though computationally tractable, SAA faces a fundamental trade-off: insufficient scenarios may fail to accurately represent the random phenomena, while larger samples increase computational demands. To alleviate this limitation, scenario reduction techniques [heitsch_scenario_2003] are often applied, but these risk omitting critical scenarios, undermining solution robustness. These shortcomings have prompted the adoption of specialised decomposition approaches, including Benders-based [sherali_modification_2002] and Lagrangian decomposition-based methods [caroe_dual_1999], which decompose the problem into smaller, more manageable subproblems. By exploiting problem structure (e.g., separability and duality), these methods mitigate SAA’s scalability challenges and reduce dependency on scenario sampling [kucukyavuz_introduction_2017]. However, their efficacy remains contingent on problem-specific properties, such as linearity or convexity, and often requires meticulous algorithmic tuning. Despite developments in decomposition methods [varga_speeding_2024], these limitations persist, driving exploration of alternative approaches that leverage machine learning (ML).
@@
Recent advances in learning-based methods for 2SP have explored surrogate models to approximate complex value functions. Several studies have proposed neural network (NN)-based approaches for learning linear programming (LP) value functions, aiming to efficiently solve two- and multi-stage stochastic programming problems [dai_neural_2021, bae_deep_2023, larsen_fast_2024, anh-nguyen_learning_2024]. Other works have focused on learning-based scenario reduction to create smaller representative scenario sets [bengio_learning-based_2020, wu_learning_2021, wu_hgcn2sp_2024]. Another line of research decomposes 2SP into multi-agent structures trained on sampled scenarios via deep reinforcement learning, with the training phase emerging as the primary computational bottleneck [hubbs_or-gym_2020, yilmaz_deep_2024, stranieri_combining_2024]. While these approaches demonstrate the advantages of surrogate modelling, they are relatively less generalisable, constrained by specific problem components or reliance on bespoke architectures. In contrast, Neur2SP (Neural Two-Stage Stochastic Programming) [dumouchelle_neur2sp_2022] presents a general framework by employing NNs to approximate the expected recourse function, motivated by the embedding of a trained NN surrogate [fischetti_deep_2018]. kronqvist_alternating_2024 further improved Neur2SP through dynamic sampling and retraining to refine the surrogate in regions of interest, a principle also embodied in the adaptive surrogate optimisation framework of liu_simulator-based_2025. However, Neur2SP’s generality comes at a cost: embedding NNs with rectified linear unit (ReLU) activations into optimisation models requires mixed-integer programming (MIP) formulations that scale with the network size, leading to computational intractability for deeper or wider architectures. This tension between generality and efficiency underscores the need for architectures that balance expressivity with tractability.
@@
Input Convex Neural Networks (ICNNs) [amos_input_2017] have emerged as a promising alternative for surrogate modelling due to their ability to fit general convex functions while retaining the flexibility and expressivity of NNs. Unlike typical feed-forward networks, ICNNs are architecturally constrained to ensure convexity with respect to their inputs, enabling exact inference through LP formulations [duchesne_supervised_2021]. This property renders them particularly advantageous for embedding within large-scale optimisation frameworks. Although ICNNs cannot model non-convex recourse functions, a limitation compared to ReLU NNs, many real-world 2SP problems in inventory management, energy dispatch, and portfolio optimisation inherently exhibit convex structure [agrawal_learning_2019, zhong_real-time_2021, luxenberg_portfolio_2024]. Despite their promise, prior research has predominantly focused on applying ICNNs for function approximation and system modelling, such as learning Optimal Power Flow (OPF) value functions [chen_data-driven_2020, rosemberg_learning_2024] or representing system dynamics in optimal control [yang_optimization-based_2021, bunning_input_2021, jiang_path-following_2022, wang_fast_2024]. However, their potential in surrogate-based optimisation, specifically for approximating convex recourse functions in 2SP, remains underexplored. As such, our work bridges this gap by unlocking efficient and scalable solutions for 2SP, leveraging the inherent convexity and computational tractability of LP.
@@
In particular, we present several interconnected contributions within a surrogate-based optimisation framework by integrating ICNNs into the 2SP paradigm, as we illustrate in Figure 1. First, we exploit the LP representability of ICNNs to eliminate integer variables inherent to conventional MIP-based embeddings. Second, we propose the first integration of ICNNs in 2SP, training them to approximate the expected value function while preserving convexity. Third, extensive numerical experiments show that our method achieves solution quality comparable to Neur2SP while reducing computational time, especially for large-scale instances, enabling application to time-critical decision problems.
@@ 2 Preliminaries
This section introduces the general formulation of 2SP and the properties of ICNNs, which form the foundation of our method.
@@ 2.1 Two-stage stochastic programming
Let $x\in\mathbb{R}^{n}$ denote first-stage decisions, $y\in\mathbb{R}^{n_{y}}$ the recourse (second-stage) decisions, and $\xi$ the relevant random variable. The stochastic components of the second-stage data $q(\xi)\in\mathbb{R}^{n_{y}}$, $W(\xi)\in\mathbb{R}^{m_{y}\times n_{y}}$, $h(\xi)\in\mathbb{R}^{m_{y}}$ and $T(\xi)\in\mathbb{R}^{m_{y}\times n}$ are determined by the realisation of $\xi$. The general linear 2SP formulation is
$\underset{\displaystyle x}{\mathrm{min}}\quad c^{\intercal}x+\mathcal{Q}(x)$ (1a)
@@
$x\geq 0,$ (1c)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$, $\mathcal{Q}(x)=\mathbb{E}_{\xi}[Q(x,\xi)]$ and the second-stage problem is
@@
which evaluates to the optimal second-stage cost under realisation $\xi$ given the first-stage decision $x$. We assume relatively complete recourse, ensuring $Q$ is feasible for any $x$ satisfying (1b) and (1c).
The decision process follows $x\rightarrow\xi\rightarrow y(x,\xi)$, where $y$ is chosen to minimise $Q(x,\xi)$. We assume $\xi$ has a known probability distribution with support $\Xi$. In practical implementations, $\Xi$ is treated as a discrete and finite set where each realisation $\xi_{s}\in\Xi$ for $s\in S\equiv\{1,\ldots,\lvert\Xi\rvert\}$ represents a scenario. This yields the deterministic equivalent formulation with a finite number of variables and constraints:
$\underset{\displaystyle x}{\mathrm{min}}\quad c^{\intercal}x+\sum_{s\in S}p_{s}q_{s}^{\intercal}y_{s}$ (3a)
@@
$T_{s}x+W_{s}y_{s}=h_{s},\quad$ $\forall s\in S$ (3c)
$y_{s}\geq 0,$ $\forall s\in S,$ (3d)
where $p_{s}$ is the probability associated with scenario $s$. This so-called extensive form (EF) (3) provides a tractable formulation. However, it suffers from scalability issues as both the number of variables and constraints increase linearly with the cardinality of the scenario set $\lvert\Xi\rvert$. As such, for variants involving mixed-integer decision variables, the computational demand increases substantially for larger scenario sets.
@@ 2.2 ICNN architectures
The surrogate function $\hat{f}$ is guaranteed to be convex in $z_{0}$ provided that $W_{1:K-1}$ are non-negative and the activation function $\sigma(\cdot)$ (ReLU in our case) is convex and non-decreasing [amos_input_2017]. This follows standard results relating to convexity-preserving operations applied to convex functions: (i) non-negative sums of convex functions are convex; (ii) the composition of a convex function with a convex non-decreasing function retains convexity. Another relevant feature of ICNNs is the presence skip connections parameterised by $S_{k}$, which directly propagates the input $z_{0}$ to deeper layers. This design aims to mitigate the restrictive effects of the non-negativity constraints imposed on $W_{k}$. Notably, these skip connections preserve convexity as the sum of a convex function (from previous layers) and a linear function (from the skip connections) remains convex.
@@ 2.3 Exact inference in ICNNs
Unlike standard NNs, which generate predictions solely through feedforward propagation of inputs across layers, the original formulation of ICNNs requires the solution of an optimisation problem as a subroutine to derive predictions. For a fully connected ICNN as described in (4), inference for a given input $z_{0}$ is mathematically equivalent to solving the following LP problem [amos_input_2017]:
$\underset{\displaystyle z_{k}}{\mathrm{min}}\quad z_{K}$ (5a)
@@
$z_{k}\geq 0,\quad$ $k=1,\ldots,K-1.$ (5c)
In this formulation, the non-linearity of ReLU activations and layer-wise equality constraints are transformed into inequalities via an epigraph representation, exploiting the fact that ReLU is a piecewise-linear function. One may note that the minimisation objective enforces that at least one constraint in (5b) or (5c) becomes active at the optimal solution, recovering the original network’s feedforward behaviour. This transformation provides a principled framework for embedding a trained ICNN surrogate into larger optimisation problems, as it offers not only an exact representation of the NN but also preserves computational tractability inherent to LP.
@@ 3 Convexity analysis
This section establishes the theoretical foundation for using ICNNs as surrogates for the second-stage cost $\mathcal{Q}(x)$ in 2SP. Since ICNNs inherently preserve convexity, their effectiveness hinges on whether $\mathcal{Q}(x)$ is convex, or whether it can be, to a sufficient degree, approximated by a piecewise-linear convex function. As $\mathcal{Q}(x)$ is evaluated by solving the second-stage problem $Q(x,\xi)$ for each scenario, it becomes a convex function of $x$ only if the second-stage problem itself is convex. We analyse three variants of 2SP problems to demonstrate that $\mathcal{Q}(x)$ retains its convexity under broad conditions.
3.1 Case 1: continuous first- and second-stage variables
When $x\in\mathbb{R}^{n}_{+}$ and $y_{s}\in\mathbb{R}^{n_{y}}_{+}$, $\forall s\in S$, the second-stage problem $Q(x,\xi_{s})$ is an LP. By LP duality, we have that
@@ 3.1 Case 1: continuous first- and second-stage variables
Each term $\pi^{\intercal}_{k}T_{s}x+\pi^{\intercal}_{k}h_{s}$ is affine in $x$, so $Q(x,\xi_{s})$ is the pointwise maximum of affine functions. As affine functions are convex and the pointwise maximum preserves convexity [boyd_convex_2004, Section 3.2.3], $Q(x,\xi_{s})$ is convex in $x$. The expected second-stage cost $\mathbb{E}_{\xi}[Q(x,\xi)]=\sum_{s\in S}p_{s}Q(x,\xi_{s})$ is a convex combination of convex functions $Q(x,\xi_{s})$, since $0\leq p_{s}\leq 1$, $\forall s\in S$, and $\sum_{s\in S}p_{s}=1$. Thus, $\mathcal{Q}(x)$ is convex in $x$.
Remark (Extension to nonlinear convex recourse).
@@
The convexity of $\mathcal{Q}(x)$ extends to second-stage problems with nonlinear convex constraints, provided that appropriate constraint qualification (CQ) conditions are satisfied, such as Slater’s CQ condition [boyd_convex_2004, Section 5.2.3], since strong duality still applies under CQ. In this more general setting, the dual formulation transforms $Q(x,\xi_{s})$ into the pointwise maximum of convex functions in $x$. Following similar duality argument, $\mathcal{Q}(x)$ preserves its convexity. Furthermore, if the first-stage problem includes a nonlinear convex objective, along with nonlinear convex inequality constraints and affine equality constraints, the recourse function remains convex in $x$ [birge_introduction_2011, Section 3.5].
3.2 Case 2: mixed-integer first- and continuous second-stage variables
Consider $x\in\mathcal{X}$ where $\mathcal{X}=\{x\in\mathbb{Z}^{n}_{+}\mid Ax=b\}$ and $y_{s}\in\mathbb{R}^{n_{y}}_{+}$. Let us define $\mathcal{X}_{\text{relax}}=\{x\in\mathbb{R}^{n}_{+}\mid Ax=b\}$ as the continuous relaxation of $\mathcal{X}$, and $\operatorname{conv}(\mathcal{X})$ as the convex hull of $\mathcal{X}$. From Case 1, we established that $\mathcal{Q}(x)$ is convex over $\mathcal{X}_{\text{relax}}$. Since $\operatorname{conv}(\mathcal{X})\subset\mathcal{X}_{\text{relax}}$, the function $\mathcal{Q}(x)$ is also convex over $\operatorname{conv}(\mathcal{X})$. By Definition 1 in [peters_convex_1986], a function $f$ defined on a possibly nonconvex set $T$ is convex if, for any finite collection $\{x^{j}\}_{j=1}^{n}\subset T$ and any convex combination $\sum_{j=1}^{n}p_{j}x^{j}\in T$, the inequality
$\sum_{j=1}^{n}p_{j}\,f(x^{j})\geq f\left(\sum_{j=1}^{n}p_{j}x^{j}\right)$ (9)
holds. Because every such convex combination of points in $\mathcal{X}$ lies in $\operatorname{conv}(\mathcal{X})$, where $\mathcal{Q}(x)$ is already known to be convex, the inequality is satisfied. Therefore, convexity follows directly for $\mathcal{X}$ once convexity over $\operatorname{conv}(\mathcal{X})$ is established. We highlight that, when $x$ is mixed-integer, the convexity of $\mathcal{Q}(x)$ trivially applies by the same reasoning.
3.3 Case 3: mixed-integer first- and second-stage variables
When both integer and continuous variables appear in the second stage, the recourse problem becomes the value function of a parametric mixed-integer linear program (MILP) in the parameter $x$. In contrast with the continuous settings of Cases 1 and 2, the value function $Q(x,\xi_{s})$ of a parametric MILP is nonconvex in general. This is well established; see hassanzadeh_value_2014 for a detailed analysis of how the feasibility of integer assignments changes with the right-hand side parameter, producing nonconvex regions in the value function. An illustrative example is provided in A.
3.3.1 A special case: quasi-convexity
Even though convexity may fail, quasi-convexity may still arise in structured settings, and ICNNs often approximate it effectively in practice. To see this, consider the sublevel set
$L_{\alpha}=\{x\mid Q(x,\xi_{s})\leq\alpha\},$ (10)
which are required to be convex for $Q(x,\xi_{s})$ to be quasi-convex. In the mixed-integer setting, the recourse function for a fixed scenario can be written as
$Q(x,\xi_{s})=\min_{y^{\mathrm{int}}\in\mathcal{Y}}\;Q_{y^{\mathrm{int}}}(x,\xi_{s}),$ (11)
where $y^{\mathrm{int}}$ denotes the second-stage integer variables and $\mathcal{Y}$ is the finite feasible set. For each fixed integer assignment $y^{\mathrm{int}}$, the remaining second-stage problem reduces to a continuous LP in the continuous recourse variables. The corresponding value function $Q_{y^{\mathrm{int}}}(x,\xi_{s})$ is convex in $x$, as analysed in Cases 1 and 2.
Hence, the sublevel set
$L_{\alpha}=\bigcup_{y^{\mathrm{int}}\in\mathcal{Y}}\{x\mid Q_{y^{\mathrm{int}}}(x,\xi_{s})\leq\alpha\}.$ (12)
is a union of convex sets. Thus, $Q(x,\xi_{s})$ is quasi-convex whenever the union is convex, which occurs for example when:
  1. (i)
    the sets $\{x\mid Q_{y^{\mathrm{int}}}(x,\xi_{s})\leq\alpha\}$ are nested as $y^{\mathrm{int}}$ varies, or
  2. (ii)
    the feasible regions associated with different integer assignments produce LP subproblems whose sublevel sets overlap in a way that yields a convex union, or
  3. (iii)
    integer patterns activate in a monotone or threshold-dependent manner with respect to $x$, so that switching between convex regimes does not introduce nonconvexity in the union.
Scenario-wise quasi-convexity does not guarantee quasi-convexity of $\mathcal{Q}(x)$ since quasi-convexity is not preserved under weighted summation. However, when the scenario-wise functions share similar convex regimes or threshold structure, the expected recourse may exhibit approximate quasi-convexity or near convexity in practice. In such settings, ICNNs are well-suited to approximate such behaviour because the convexity of their output in $x$ captures the dominant convex shape of each regime while smoothing over the transitions between regimes.
3.4 Practical implications for ICNN surrogates